OpenStudy (matthew120):

State whether the ratios form a proportion. 25/35 and 35/45 A. Yes B. No

OpenStudy (smilie2015):

no(B) 25/35 = 35/45 5/7 = 7/9 5*9=7^2 45=49

OpenStudy (smilie2015):

3/9=p/24 by cross multiplication 3*24=9p 72=9p dividing both sides with 9 8=p

OpenStudy (matthew120):

@mathmale

OpenStudy (smilie2015):

g/6=3/36 it will use the same method as mentioned above 36g=18 g=18/36 g=1/2 answer B

OpenStudy (matthew120):

@imqwerty

OpenStudy (smilie2015):

for the above one boys:girls 3:4=x:16 3/4=x/16 i think u can solve it further

OpenStudy (matthew120):

ok i need no more help